given two points P(2,1) and Q(4,7).find an equation of the straight line passing through Q and perpendicular to PQ
1
Expert's answer
2020-11-04T16:41:17-0500

Slope

"m = \frac{y_2-y_1}{x_2-x_1} = \frac{7-1}{4-2} = 3"

Perpendicular slope

"m = -\frac{1}{3}"

Passes through (4,7)

"y - y_1 = m(x - x_1)"

"y - 7 = -\frac{1}{3}(x - 4)"

"y = -\frac{1}{3}x + 7 + \frac{4}{3}"

"y = -\frac{1}{3}x + \frac{25}{3}"